3. Evaluate the limit

lim(x,y)->(2,2) (x-y)/(x^4-y^4)
1
Expert's answer
2020-09-29T18:18:18-0400

"x^4-y^4=(x^2)^2-(y^2)^2"

Using the formula "(a^2-b^2)=(a+b)(a-b)," we get

"x^4-y^4=(x^2+y^2)(x^2-y^2)"

Again, "(x^2-y^2)" can be written as "(x+y)(x-y)"

"\therefore x^4-y^4=(x^2+y^2)(x+y)(x-y)"

Now "\lim_{(x,y)\rightarrow(2,2)}\frac{x-y}{x^4-y^4}"

"=\lim_{(x,y)\rightarrow(2,2)}\frac{x-y}{(x^2+y^2)(x+y)(x-y)}"

"=\lim_{(x,y)\rightarrow(2,2)}\frac{1}{(x^2+y^2)(x+y)}"

"=\frac{1}{(2^2+2^2)(2+2)}"

"=\frac{1}{32}"

"\therefore \lim_{(x,y)\rightarrow(2,2)}\frac{1}{(x^2+y^2)(x+y)}=\frac{1}{32}"